OpenStudy (anonymous):

-1/9(x-27)+1/3(x+3)=x-4 It says to solve the equation and check the solution. Any help would be greatly appreciated

OpenStudy (anonymous):

this is how i am interpreting your question i hope it is right. \[-\frac{ 1 }{ 9 }\left( x-27 \right)+\frac{ 1 }{ 3 }\left( x+3 \right)=x-4\]

OpenStudy (anonymous):

multiply through by 9 -(x-27)+3(x+3)=9x-36 open brackets -x+27+3x+9=9x-36 collect like terms 7x=72 simplify (i'll leave that to you

OpenStudy (anonymous):

lemme know if u need more help